OpenStudy (anonymous):

Select the inequality that models the problem. The length of a rectangle is twice its width. If the perimeter of the rectangle is less than 50 meters, what is the greatest width of the rectangle?

OpenStudy (anonymous):

2 • 2w + 2w < 50 2 + 2w < 50 2w • w < 50 2w < 50

OpenStudy (anonymous):

option a

OpenStudy (anonymous):

the reason being that...if w is the width of the rectangle, then 2w will be the length of the rectangle....... perimeter is = 2(l+b) =2(2w+w)=2.2w+2w....got it?

OpenStudy (anonymous):

this 2.2w+2w has to be less that 50

OpenStudy (trojanpoem):

L = 2 W 2L + 2W < 50 4W + 2W < 50 6W < 50 W < 25/3